OpenStudy (anonymous):

The product of the roots of ax^2 + bx + (a^2 + 1) =0 is the sum of the roots of x^2 - (2a + 1) x + b =0 , Then a+1 =?

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

product of the roots =c/a =(a^2 + 1)/a sum of the roots=-b/a =(2a + 1)/1 product of the roots of = the sum of the root (a^2 + 1)/a=(2a + 1) u proceed nxt

OpenStudy (anonymous):

i got it thxxx...))